OpenStudy (anonymous):

The table shows the heights in inches of trees after they have been planted. Determine the equation which relates x and y. Height In Pot, x | Height Without Pot, y 30 | 18 36 | 24 42 | 30 48 | 36 A. y = x - 6 B. y = x - 10 C. y = x - 12 D. y = x + 6

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

30=18+12 hence 18=30-12 ?

OpenStudy (anonymous):

I'm....confused. Sorry........@surjithayer

OpenStudy (anonymous):

30=18+12 x=y+12 x-12=y

OpenStudy (anonymous):

Here...I'll tag someone that usually will help you. @texaschic101

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

Welcome

OpenStudy (texaschic101):

take 2 sets of points and find the slope (30,18) and (36,24) now plug that into the slope formula slope(m) = (y2 - y1) / (x2 - x1) (30,18)....x1 = 30 and y1 = 18 (36,24)...x2 = 36 and y2 = 24 now sub slope(m) = (24 - 18) / (36 - 30) slope(m) = 6/6 which equals 1 now we take y = mx + b slope(m) = 1 you can use either of your points...(30,18)...x = 30 and y = 18 now we sub y = mx + b 18 = 1(30) + b 18 = 30 + b 18 - 30 = b -12 = b so your equation is : y = 1x - 12 or just y = x - 12

OpenStudy (anonymous):

Thank you!!!! :)